OpenStudy (anonymous):

A three-character password consists of 2 different digits between 0 and 9, inclusive, and 1 letter of the English alphabet. The letter must appear as first or second character. How many different passwords are possible?

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

there may be three two cases i hope 1st alphabet and other two numbers 2nd alphabet and 1st and last numbers in the first case: 26p1 + 10p2= 26+ 10*9= 116 and the second case is 10P1 + 26p1 + 10p1 = 46 so the sum will be the answer I hope...